OpenStudy (lisa2345):

Use the given intercepts to determine the correct equation of the line. 9x + y = 20 5x + 4y = 9 4x + 5y = 20 4x - 5y = 9 y-intercept (0,4) x-intercept (5,0) plz just give answer I need it right away

OpenStudy (anonymous):

find the slope first by using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (0,4)..x = 0 and y = 4 (5,0)..x = 5 and y = 0 now we sub slope(m) = (0 - 4) / (5 - 0) slope(m) = -4/5 now we use y = mx + b slope(m) = -4/5 use either of your sets of points...(5,0)..x = 5 and y = 0 now sub 0 = -4/5(5) + b 0 = -4 + b 4 = b so the equation is : y = -4/5x + 4 now we will put it in standard form Ax + By = C y = -4/5x + 4 --- multiply everything by 5 to get rid of the fraction 5y = -4x + 20 -- add 4x to both sides 4x + 5y = 20 <== standard form equation

OpenStudy (anonymous):

@kelliegirl33 i need help

OpenStudy (lisa2345):

get off my question you can message her >:

OpenStudy (anonymous):

ohh who asked YOU

OpenStudy (anonymous):

I will be with you in a second lala...hold on. lisa, do you have any questions ?

OpenStudy (anonymous):

ok

OpenStudy (lisa2345):

Nobody asked dicchead , , and yess I do

OpenStudy (anonymous):

watch your language

OpenStudy (anonymous):

DICCHHEAD IS RIGHT HERE

OpenStudy (anonymous):

whats your question lisa

OpenStudy (lisa2345):

@kelliegirl133 I don't need your help if you're gonna tell me to watch my language like im a little retricekid . I am a junior not a 6th grader

OpenStudy (anonymous):

doesn't matter what age you are...there are younger kids on this site and you are not setting a very good example.

OpenStudy (anonymous):

I understand..I get upset too and may say some bad things....but I can control my language on a site such as this. Thats just being mature

OpenStudy (anonymous):

now if you want, I can answer some questions that you might not understand ? Its up to you ?

OpenStudy (lisa2345):

I don't care you disrespect you get disrespected point blank period .

OpenStudy (anonymous):

I admit Lisa should not have jumped in on your question to ask a question, but there was no need for that language. You could have just mentioned to her that it was not cool to do that

OpenStudy (anonymous):

oops...lala

OpenStudy (lisa2345):

no thanks hop off my question boo (peace) Don't need your lecture

OpenStudy (anonymous):

i didint do anything i just asked for help to you it was none of her bussiness

OpenStudy (anonymous):

good luck figuring out your math :)